OpenStudy (anonymous):

what is the factorization of 100x^2+20x+1

OpenStudy (anonymous):

would it (2x+10)(2x+10)

OpenStudy (anonymous):

no but it is another perfect square

OpenStudy (anonymous):

what is \(100x^2\) the square of ?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

10

OpenStudy (anonymous):

actually \(10x\) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok, and what is 1 the square of?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now we check that \(2\times 10\times 1=20\) which is right so we have the perfect square \((10x+1)^2\)

OpenStudy (anonymous):

would just x+10 x+10 right

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

\(100x^2\) is the same as \((10x)^2\) so \(10x\) will be the first term

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and \(1^2=1\) so \(1\) will be the second term

OpenStudy (anonymous):

yeah i just wrong them wrong

OpenStudy (anonymous):

since \(2\times 10x\times 1=20x\) which is the middle term, we know it is a perfect square

OpenStudy (anonymous):

the square is \((10x+1)^2\)